Here's a question for you to solve and it has given me trouble. Here'a the problem 1/2k-3=2-3/4k
Answer by i dont like math(2)

all you have to do is isolate K
1/2k-3=2-3/4k
-3/4k -3/4k
-1/4k-3=2
+3 +3
-1/4k=5
________
-1/4 -1/4
k=-20

Hope that helps